The perimeter of a regular polygon is 48 cm and its apothem is 6 cm. What is the area of the polygon?


1
Expert's answer
2021-01-07T15:59:59-0500

The perimeter of a regular polygon is 48cm => P = 48cm and apothem is 6 cm => r = 6cm;


The formula of Regular polygon Circumscribing a circle is:

"Perimeter = 2nrtan(\frac{180^o}{n})" and "Area = nr^2tan(\frac{180^o}{n})"

In the perimeter formula, see that "tan(\frac{180^o}{n}) = \frac{Peimeter}{2 * nr}," substitute it to the Area:

"Area = nr^2*(\frac{P}{2*nr}) = \frac{1}{2}Pr = \frac{1}{2}*48*6 = 144cm^2"

Answer, Area = 144cm2